OpenStudy (anonymous):

7x - 3y = 4 2x - 4y = 1 The solution to the system of equations is _____. (Be sure to show all work.)

OpenStudy (anonymous):

We have done several of these, what is the first step?

OpenStudy (anonymous):

5x -1y =6 am i right meverett

OpenStudy (anonymous):

i havent done one that are subtration

OpenStudy (anonymous):

@sunset almost Take 7x - 3y = 4 2x - 4y = 1 We need to make the y's opposite so multiple the first equation by 4 28x - 12y = 16 and the second equation by 3 6x - 12y = 3 and take these two equations 28x - 12y = 16 6x - 12y = 3 and subtraction 22x = 13 and so on...